OpenStudy (anonymous):

A train leaves a station at the 0.0 m distance marker travelling with a constant velocity of 36.o m/s. a) How many seconds later will the train pass the 1620 m marker? b) what is the velocity of the train in km/h? (129.6 km/hr)

OpenStudy (shamim):

we know the formula time=distance/velocity=1620/36

OpenStudy (shamim):

36m/s=(36*3600/1000) km/ h

OpenStudy (shamim):

because u know the formula 1 km=1000m 1 h=3600 sec